OpenStudy (anonymous):

14. In rhombus LMNO, the ration of Lm to MN is 4:3. Find LM if the perimeter is 42 inches.

OpenStudy (anonymous):

ratio*

OpenStudy (asnaseer):

lets call length of LM 'a', and length of MN 'b'. then we know:\[\frac{a}{b}=\frac{4}{3}\]\[2a+2b=42\]divide both sides by 2 to get:\[a+b=21\]so:\[b=21-a\]put this into 1st equation to get:\[\frac{a}{21-a}=\frac{4}{3}\]cross multiply to get:\[3a=84-4a\] you should be able to solve from here....

OpenStudy (asnaseer):

I hope you understood the steps taken above...